### v1

# Nonparametric Estimation of Joint Entropy through Partitioned Sample-Spacing Method

Abstract
We propose a nonparametric estimator of multivariate joint entropy based on partitioned sample spacings (PSS). The method extends univariate spacing ideas to $\mathbb{R}^{d}$ by partitioning into localized cells and aggregating within-cell statistics, with strong consistency guarantees under mild conditions. In benchmarks across diverse distributions, PSS consistently outperforms $k$-nearest neighbor estimators and achieves accuracy competitive with recent normalizing flow–based methods, while requiring no training or auxiliary density modeling. The estimator scales favorably in moderately high dimensions (d=10d=104040) and shows particular robustness to correlated or skewed distributions. These properties positions PSS as a practical alternative to NF-based approaches, with broad potential in information-theoretic machine learning applications.
00footnotetext: Corresponding author: hoo0321@yonsei.ac.kr 1Department of Statistics and Data Science, Yonsei University 2Department of Statistics and Probability, Michigan State University
1 Introduction
Entropy and mutual information are fundamental quantities in data analysis and machine learning, as they quantify uncertainty and dependence between variables Shannon1948. Unlike correlation, which captures only linear relationships, they can detect general nonlinear dependencies and remain invariant under monotone transformations. These properties have made entropy and mutual information central tools for feature selection, representation learning, and causal discovery.
Estimating entropy and mutual information from samples is challenging. Among nonparametric approaches, the $k$-nearest neighbor (kNN) estimator is widely used PhysRevE.69.066138; Paninski20031191. It requires no distributional assumptions and adapts naturally to new data, but its accuracy degrades in multivariate settings, particularly under strong correlation or high dimension. More broadly, there is a long line of work on nonparametric estimation of divergences and conditional information using kkNN statistics poczos, which provides consistency guarantees but suffers from the same scalability issues.Recent refinements such as bias-corrected weighting BerrettSamworthYuan2019 improve asymptotic properties, yet all kNN variants rely on costly neighbor searches and remain sensitive to dependence.
Another line of work leverages copula decompositions. Ariel and Louzoun copula proposed the Copula Decomposition Entropy Estimator (CADEE), which separates marginal and copula contributions using Sklar’s theorem. By recursively estimating copula entropy on the compact $0,1]^{d}$ domain, CADEE achieves strong robustness and scales better than kNN, particularly when supports are mixed or unbounded.
In parallel, trainable normalizing flow methods have emerged. Ao and Li AoLi2022 introduced an estimator that learns a bijection to a tractable base distribution, while Butakov and Tolmachev ButakovTolmachev2024 proposed MIENF, which fits a pair of flows with finite-sample guarantees. Saad et al. saad developed inference-based entropy estimators (EEVI) that compute bounds within probabilistic generative models. While such estimators can reduce bias, they require GPU-scale training or access to tractable model densities, and may fail when the model family is misspecified.
In this work, we revisit Vasicek’s classical sample-spacing idea Dudewicz1986161 and extend it to the multivariate setting by applying it locally within partitions of the sample space. The resulting estimator is fully nonparametric, avoids both neighbor search and neural training, and remains robust under heavy tails and nonlinear dependence. Under mild assumptions—boundedness and continuity of the density within each partition—we prove strong consistency, with an additional moment condition for entropy. Empirically, our method consistently outperforms kNN in mean-squared error, improves upon CADEE in correlated and moderately high-dimensional settings, and achieves accuracy competitive with normalizing-flow approaches while being substantially more efficient.
2 Preliminary
2.1 Nearest Neighbor–Based Estimators
One widely used class of nonparametric entropy estimators relies on $k$-nearest neighbor (kNN) distances. Given $n$ i.i.d. samples $X_{1},\dots,X_{n}\in\mathbb{R}^{d}$, let $R_{i,k}$ denote the distance from $X_{i}$ to its $k$th nearest neighbor. These distances provide local volume estimates, which can be inverted to obtain density estimates and hence entropy.
Kozachenko–Leonenko (KL) estimator.
The classical estimator of Kozachenko assumes local uniformity within a ball of radius $R_{i,k}$ around $X_{i}$. With $V_{d}=\pi^{d/2}/\Gamma(1+d/2)$ the volume of the $d$-dimensional unit ball, the KL estimator is
$\hat{H}_{\mathrm{KL}}=-\psi(k)+\psi(n)+\log V_{d}+\frac{d}{n}\sum_{i=1}^{n}\log R_{i,k},$
where $\psi$ is the digamma function. Despite its simplicity, the KL estimator can suffer from large bias in high dimensions gao2015efficientestimationmutualinformation.
Kraskov–Stögbauer–Grassberger (KSG) estimator.
kraskov2004 proposed a modification using hyper-rectangles instead of balls. For each $X_{i}$, let $R_{i,j}$ be the distance to its $k$th nearest neighbor along coordinate $j$. Defining a rectangle with side lengths $\{R_{i,j}\}$, the KSG entropy estimate becomes
$\hat{H}_{\mathrm{KSG}}=-\psi(k)+\psi(n)+\frac{d-1}{k}+\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{d}\log R_{i,j}.$
Originally introduced for mutual information, KSG often improves stability compared to KL, especially under strong dependence.
Truncated estimators and Normalizing Flows.
Boundary bias can be severe when the support is compact. To address this, Ao and Li AoLi2022 introduced truncated versions (tKL, tKSG), where neighbor cells are clipped to the support $0,1]^{d}$. They proved that tKL is unbiased for uniform distributions. Building on this property, they proposed a normalizing flow transformation that maps samples approximately to uniform, and then applies the truncated estimator with a Jacobian correction. This yields the so-called uniformizing mapping (UM) estimators (UM-tKL, UM-tKSG), which achieve substantially lower bias in high dimensions. Normalizing flows themselves are neural network models of invertible transformations RezendeMohamed2015; Papamakarios2021. They require GPU-based training with many gradient updates, making them far more computationally demanding than nonparametric methods. Thus NF-based entropy estimation can reach high accuracy but at the cost of significant training overhead.
2.2 Univariate Case
Let $x_{(1)}<x_{(2)}<\cdots<x_{(n)}$ be order statistics of an i.i.d. sample $x_{1},x_{2},\dots,x_{n}$ drawn from a continuous random variable $X$ with probability density function $f(x)$. The difference $x_{(i+m)}-x_{(i)}$ is called m-spacing Learned-Miller20041271. Using this idea, park_and_park proposed the following density estimator based on sample spacings:
$\hat{f}_{n}(x)=\begin{cases}\displaystyle\frac{2m}{n\big(x_{(i+m)}-x_{(i-m)}\big)}&\text{for }\xi_{i}<x\leq\xi_{i+1}\\ 0&\text{otherwise}\end{cases}$
where $\xi_{i}=\frac{x_{(i-m)}+\cdots+x_{(i+m-1)}}{2m}$, and $x_{(i)}=x_{(n)}$ for $i>n$. Note that for $i<1$, $x_{(i)}=x_{(1)}$ and for $i>n$, $x_{(i)}=x_{(n)}$. Based on this m-spacing method differential entropy is derived as (Vasicek, 1976)
$H_{v}(n,m)=\frac{1}{n}\sum_{i=1}^{n}\log\left(\frac{n}{2m}(x_{(i+m)}-x_{(i-m)})\right)$
This entropy estimator converges to $H(f)$ as $n,m\to\infty$ and $\frac{m}{n}\to 0$ Dudewicz1986161.
While Vasicek proved the consistency of the entropy estimator $H_{v}(n,m)$, he did not establish the consistency of the corresponding density estimator. We fill this gap by establishing the almost sure consistency of $\hat{f_{n}}(x)$, under mild assumptions on the spacing parameter $m$. The key step is to prove that the averaged grid point $\xi_{i}$ converges to the order statistic $X_{(i)}$, ensuring that the estimator localizes correctly around the sample point.
Lemma 1 (Consistency of $\xi_{i}$).
If $m\rightarrow\infty$ and $m/n\rightarrow 0$ as $n\rightarrow\infty$, then $\xi_{i}-X_{(i)}$ converges to 0 almost surely.
The choice of the spacing parameter $m$ is critical for balancing bias and variance. To ensure consistency, $m$ must grow with the sample size nnBert-vanEs. For our theoretical results and empirical studies, we adopt the widely used rate $m=\lfloor\sqrt{n}+1/2\rfloor$ Crzc, which is known to be effective for smooth distributions.
Theorem 1 (Consistency of $\hat{f}_{n}(x)$).
If the spacing parameter is set as $m=\lfloor\sqrt{n}+1/2\rfloor$, then the sample-spacing density estimator $\hat{f}_{n}(x)$ converges almost surely to $f(x)$.
$\hat{f}_{n}(x)\xrightarrow{a.s.}f(x)$
where $\xrightarrow{a.s.}$ denotes almost sure convergence.
This result formally establishes the almost sure consistency of the univariate sample-spacing density estimator, providing a theoretical foundation for its extension to the multivariate and partitioned settings developed in the next sections.
3 Proposed Extension
3.1 Bivariate Case
Let $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ be i.i.d. samples from a continuous joint density $f(x,y)$ defined on $\mathbb{R}^{2}$. The joint (differential) entropy is
$H(X,Y)=-\iint_{\mathbb{R}^{2}}f(x,y)\log f(x,y)\,dx\,dy,$
which measures the overall uncertainty in the bivariate distribution.
Vasicek’s sample-spacing method was originally proposed for univariate entropy estimation. However, the same idea can be extended to higher dimensions by constructing a piecewise-constant joint density estimator on a grid formed by the order statistics of each variable. This generalization retains the nonparametric nature of the method while enabling it to capture relationships between multiple variables.
Let $x_{(1)}\leq\dots\leq x_{(n)}$ and $y_{(1)}\leq\dots\leq y_{(n)}$ be the order statistics of $\{X_{i}\}$ and $\{Y_{i}\}$, respectively. Define the averaged grid points $\xi_{i}=\frac{x_{(i-m_{x})}+\cdots+x_{(i+m_{x}-1)}}{2m_{x}}$, $\eta_{j}=\frac{y_{(j-m_{y})}+\cdots+y_{(j+m_{y}-1)}}{2m_{y}}$.
with $x_{(i)}=x_{(1)}$ for $i<1$, $x_{(i)}=x_{(n)}$ for $i>n$, and analogously for $y_{(j)}$.
If $X$ and $Y$ are independent, the joint density factorizes as $f(x,y)=f_{X}(x)f_{Y}(y)$. In this case, the joint density estimator can be constructed as the product of two univariate spacing-based estimators $\hat{f}_{n}^{\mathrm{ind}}(x,y)$:
$\frac{4m_{x}m_{y}}{n^{2}\,(x_{(i+m_{x})}-x_{(i-m_{x})})\,(y_{(j+m_{y})}-y_{(j-m_{y})})}$
for $\xi_{i}<x\leq\xi_{i+1},\ \eta_{j}<y\leq\eta_{j+1}$, $i,j=1,\dots,n$. The corresponding joint entropy estimator is then
$H_{v}^{\mathrm{ind}}(n,m_{x},m_{y})=-\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\log\hat{f}_{n}^{\mathrm{ind}}(x,y).$
A direct extension of the univariate sample-spacing method to the bivariate case constructs a grid from marginal order statistics. This is simple but flawed: marginal grids ignore joint geometry, so dense regions may be covered by large low-density cells and vice versa. For correlated data, the estimator underestimates along dense diagonals and overestimates in sparse regions, because marginal orderings distort the grid.
Figure 1(a) shows this failure for a bivariate normal with $\rho=0.8$. The direct method does not reflect the linear dependence and even assigns spurious high density to off-diagonal regions. To address this, we partition each axis into $\ell$ equal-width segments and apply the sample-spacing estimator locally. This prevents mis-assignment across quadrants and better captures the correlation, as seen in Figure 1(b) with $\ell=3$.
Unlike adaptive partitioning approaches based on independence tests Darbellay19991315, or sample-spacing methods assuming radial symmetry Lee20102208, our method uses fixed, assumption-free partitions, making it both simple and broadly applicable.
Figure 1: Comparison of density estimation for a bivariate normal with $\rho=0.8$. The direct method (left) fails to capture the correlation, while the partitioned method (right) successfully reflects the joint structure.
3.2 The Partitioned Density Estimator
Let $P_{k}$ be the $k$-th partition, $m_{k}$ be the spacing parameter in $P_{k}$, $n_{k}$ be the number of samples in that partition, and $(x_{(1)}^{k},\dots,x_{(n_{k})}^{k})$, $(y_{(1)}^{k},\dots,y_{(n_{k})}^{k})$ be the corresponding order statistics for $k=1,\cdots,\ell^{2}$. To ensure the estimator integrates to 1, we weigh the local density estimate in the $k$-th partition by the sample proportion $\frac{n_{k}}{n}$. The partitioned sample-spacing density estimator $\hat{f}_{n,\ell}(x,y)$ is:
$\frac{n_{k}}{n}\frac{4m_{k}^{2}}{n_{k}^{2}(x_{(i+m_{k})}^{k}-x_{(i-m_{k})}^{k})(y_{(j+m_{k})}^{k}-y_{(j-m_{k})}^{k})}$
where $(x,y)\in(\xi_{i}^{k},\xi_{i+1}^{k}]\times(\eta_{j}^{k},\eta_{j+1}^{k}]$. The grid points are defined as $\xi_{i}^{k}=\frac{x_{(i-m_{k})}^{k}+\dots+x_{(i+m_{k}-1)}^{k}}{2m_{k}}$ and $\eta_{j}^{k}=\frac{y_{(j-m_{k})}^{k}+\dots+y_{(j+m_{k}-1)}^{k}}{2m_{k}}$, with boundary points $\xi_{0}^{k}=x_{(1)}^{k}$ and $\xi_{n_{k}+1}^{k}=x_{(n_{k})}^{k}$. Analogously, $\eta_{0}^{k}=y_{(1)}^{k}$ and $\eta_{n_{k}+1}^{k}=y_{(n_{k})}^{k}$. The partitioned density estimator is defined locally within each partition. To make the estimator more adaptive and reduce estimation bias, the sub-grid is constructed on the empirical support of the data within each partition, i.e., on the partition $\xi_{0}^{k},\xi_{n_{k}+1}^{k}]\times[\eta_{0}^{k},\eta_{n_{k}+1}^{k}]$. While this means the density is implicitly zero in the regions between the data’s range and the partition boundaries for finite samples, this choice is empirically superior, yielding a significantly lower MSE. As a result of this data-driven construction, the estimator is not guaranteed to integrate to exactly 1 for finite samples; however, it is asymptotically integrates to 1, as the empty regions between the data’s range and the partition boundaries vanish as $n_{k}\rightarrow\infty$.
Proposition 1.
The partitioned sample-spacing estimator $\hat{f}_{n,\ell}(x,y)$ is asymptotically a valid probability density function, satisfying
$\lim_{n\to\infty}\int_{\mathbb{R}^{2}}\hat{f}_{n,\ell}(x,y)\,dx\,dy=1$
This estimator is a valid probability density function and is strongly consistent. We formalize this in the following lemma and theorem. The proof strategy is to first establish, in lemma 2, the almost sure convergence of the univariate $m$-spacing estimators for the conditional distributions within each partition. This result serves as the core ingredient for Theorem 2, which combines these components to prove the consistency of the overall joint density estimator.
The assumptions in the lemma are critical for this two-stage argument. The conditions on $\ell(n)$ and $n_{k}$ ensure that as the total sample size $n$ grows, the partitions shrink appropriately while still accumulating enough data for asymptotic analysis. Similarly, the condition on the spacing parameter $m_{k}$ is crucial. As established in the preceding section, we adopt the specific rate $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$, which is required to prove convergence. Hence, we assume the conditions on $\ell(n)$, $n_{k}$, and the specific rate for $m_{k}$ holds for all subsequent asymptotic results.
Lemma 2.
Let the event $C_{k}\;=\;\bigl\{(X,Y)\in P_{k}\bigr\},$ $n_{k};=\sum_{v=1}^{n}\mathbf{1}\{(X_{v},Y_{v})\in C_{k}\}.$ Assume that as $n\to\infty$, $\ell(n)\to\infty,\ \ell(n)^{2}=o(n),\ n_{k}\rightarrow\infty.$ Further, assume the spacing parameter $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$. For $x\in(\xi_{j}^{k},\xi_{j+1}^{k}],\ y\in(\eta_{i}^{k},\eta_{i+1}^{k}]$ define the univariate mm‐spacing estimators
$\widehat{f}_{X,k}(x)=\frac{2\,m_{k}}{\,n_{k}\bigl(x_{k,(i+m_{k})}-x_{k,(i-m_{k})}\bigr)}$
$\widehat{f}_{Y,k}(y)=\frac{2\,m_{k}}{\,n_{k}\bigl(y_{k,(j+m_{k})}-y_{k,(j-m_{k})}\bigr)}.$
Then for any fixed $(x,y)\in(\eta_{i}^{k},\eta_{i+1}^{k}]\times(\xi_{j}^{k},\xi_{j+1}^{k}]$,
$\widehat{f}_{X,k}(x)\;\xrightarrow{a.s.}\;f_{X\mid C_{k}}(x),\ \widehat{f}_{Y,k}(y)\;\xrightarrow{a.s.}\;f_{Y\mid C_{k}}(y).$
Theorem 2 (Consistency of $\hat{f}_{n,\ell}(x,y)$).
The partitioned bivariate density estimator
$\hat{f}_{n,\ell}(x,y)=\frac{n_{k}}{n}\cdot\hat{f}_{X;k}(x)\cdot\hat{f}_{Y;k}(y),$
converges almost surely to $f(x,y)$.
The almost sure convergence in Theorem 2 provides the foundation for $L^{1}$ convergence via Scheffé’s theorem Williams: if a sequence of nonnegative integrable functions $g_{n}$ satisfies (i) $g_{n}\to g$ a.e. and (ii) $\int g_{n}\to\int g$, then $\int|g_{n}-g|\to 0$. In our setting, Theorem 2 yields (i) for $\hat{f}_{n,\ell}$, and Proposition 1 establishes (ii) with $\int f=1$.
Corollary 1 ($L^{1}$ convergence of $\hat{f}_{n,\ell}$).
By Scheffé’s theorem, the L1 distance between the estimator and the true density converges to zero:
$\int_{\mathbb{R}^{2}}\big|\hat{f}_{n,\ell}(x,y)-f(x,y)\big|\,dx\,dy\;\longrightarrow\;0.$
3.3 The Joint Entropy Estimator
Given the consistent density estimator $\hat{f}_{n,\ell}(x,y)$, a natural and powerful way to estimate the joint entropy is via a ”plug-in” approach. We replace the true density $f$ with our estimator $\hat{f}_{n,\ell}$ in the definition of entropy and approximate the expectation using the empirical distribution of the $n$ observed samples. This leads to the final entropy estimator:
$\hat{H}_{n,\ell}(X,Y)=-\frac{1}{n}\sum_{v=1}^{n}\log\hat{f}_{n,\ell}(X_{v},Y_{v}).$
By applying the partition parameter $\ell$, several advantages arise. Since this estimator is computed locally, it is expected to perform well on highly correlated data and to effectively capture details of the distribution, such as skewness and kurtosis.
Additionally, the number of grids formed by the partitioned model is considerably smaller than in the previous model, making it more computationally efficient. For example, when $n$ samples are given, the previous model forms $n^{2}=\left(\sum_{k=1}^{\ell}n_{k}\right)^{2}$ grids, whereas the partitioned model forms $\sum_{k=1}^{\ell}n_{k}^{2}$ grids. As a result, the model becomes more computationally efficient as $\ell$ increases. However, factors such as variance, correlation, and sample size should also be considered when choosing the optimal value of $\ell$.
The optimal $\ell$ is expected to be invariant to variance, since the length of each partition is set as $\frac{\max(x)-\min(x)}{\ell}$ and $\frac{\max(y)-\min(y)}{\ell}$. Since these lengths are proportional to the variances of $X$ and $Y$, the number of samples $n_{k}$ in each partition is not influenced by variance for a fixed $\ell$. We can formally demonstrate this.
Proposition 2.
The partition parameter $\ell$ is location invariant and scale equivariant.
Building upon the consistency of the density estimator, we now establish the $L^{1}$ convergence of the plug-in entropy estimator. The crucial step is to prove the uniform integrability of the log-density estimator sequence. This requires the globally adopted rate for the spacing parameter, $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$, as well as a mild integrability condition on the true density $f$ itself. The following theorem formalizes this main result.
Theorem 3.
Assume that the true density $f(x,y)$ satisfies $\log f\in L^{1+\delta}$ for some $\delta\in(0,1)$. Let the spacing parameter be $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$ with $n_{k}\geq 1$, and let the number of partitions satisfy $\ell(n)^{2}=o(n)$. Then, as $n\to\infty$
$\hat{H}_{n,\ell}(X,Y)\xrightarrow{L^{1}}H(f):=\mathbb{E}[-\log f(X,Y)]$
4 Multivariate Extension
4.1 Multivariate Density Estimator
The partitioned sample-spacing method, initially developed for the bivariate case, can be extended to multivariate scenarios. Let $\mathbf{X}=(X_{1},X_{2},\dots,X_{d})$ represent a $d$-dimensional random vector, with $n$ samples $\{(x_{v1},x_{v2},\dots,x_{vd})\}_{v=1}^{n}$. The partition parameter $\ell$ divides each dimension into $\ell$ segments, resulting in $\ell^{d}$ hyper-rectangular partitions, where each partition has a size proportional to the range of the data along each axis, i.e., $\frac{\max(X_{i})-\min(X_{i})}{\ell}$ for $i=1,\dots,d$.
The density function estimator $f_{n,\ell}(\mathbf{X})$ is generalized to the multivariate case as follows:
$\frac{n_{k}}{n}\cdot\prod_{i=1}^{d}\frac{2m_{k}}{n_{k}(x_{i,(a_{i}+m_{k})}^{k}-x_{i,(a_{i}-m_{k})}^{k})}$
where $\xi_{a_{j}}^{k}<x_{j}\leq\xi_{a_{j}+1}^{k},j=1,\dots,d$ and $k=1,\dots,\ell^{d}$ indexes the partitions, $n_{k}$ is the number of samples in partition $k$, $m_{k}$ is the spacing parameter, $a_{i}$ is the index of the order statistic corresponding to the interval that $X_{i}$ falls into and $x_{i,(r)}^{k}$ denote the $r$th marginal order statistic for the $i$th dimension within partition $k$. The sub-grid points for each dimension is defined as $\xi_{i,a_{i}}^{k}=\frac{x_{i,(a_{i}-m_{k})}^{k}+\dots+x_{i,(a_{i}+m_{k}-1)}^{k}}{2m_{k}}$. Analogous to the bivariate case, the sub-grid is constructed on the empirical support of the data within each partition. Therefore, the outer boundaries for each dimension $j$ are set to the minimum and maximum marginal order statistics within that partition: $\xi_{j,0}^{k}=x_{j,(1)}^{k}\;\text{and}\;\xi_{j,n_{k}+1}^{k}=x_{j,(n_{k})}^{k}.$
Proposition 3.
The partitioned multivariate sample-spacing density estimator $\hat{f}_{n,l}(\mathbf{x})$ satisfies
$\lim_{n\rightarrow\infty}\int_{\mathbb{R}^{d}}\hat{f}_{n,l}(\mathbf{x})\,d\mathbf{x}=1,$
Now that we have established that $\hat{f}_{n,l}(\mathbf{x})$ is an asymptotically valid probability density function, we proceed to the main result for the estimator: proving that it is also strongly consistent. This means showing that the estimator converges almost surely to the true underlying density $f(\mathbf{x})$ as the sample size increases. The following theorem formalizes this crucial property.
Theorem 4 (Consistency of the Partitioned Multivariate Density Estimator).
The partitioned multivariate density estimator is defined as:
$\hat{f}_{n,\ell}(\mathbf{x})=\frac{n_{k}}{n}\cdot\prod_{j=1}^{d}\hat{f}_{X_{j};k}(x_{j})\text{}$
where $\hat{f}_{X_{j};k}(x_{j})=\frac{2m_{k}}{n_{k}(x_{j,(a_{j}+m_{k})}^{k}-x_{j,(a_{j}-m_{k})}^{k})}$. Assume that as $n\to\infty$, the number of partitions satisfies $\ell(n)\to\infty$ with $\ell(n)^{d}=o(n)$. Further, assume the spacing parameter $m_{k}$ satisfies the conditions $m_{k}\to\infty$ and $m_{k}/n_{k}\to 0$, which are met by the globally adopted rate of $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$. Given these conditions, $\hat{f}_{n,\ell}(\mathbf{x})$ converges almost surely to the true joint density $f(\mathbf{x})$:
$\hat{f}_{n,\ell}(\mathbf{x})\xrightarrow{a.s.}f(\mathbf{x})$
4.2 Multivariate Joint Entropy Estimator
The multivariate joint entropy estimator is defined as the plug-in estimator based on the partitioned multivariate density:
$H_{v}(\mathbf{X},n,\ell)=-\frac{1}{n}\sum_{v=1}^{n}\log\hat{f}_{n,\ell}(\mathbf{X}_{v}).$
For a given data point $\mathbf{X}_{v}$ that falls into a sub-grid cell indexed by $\mathbf{a}=(a_{1},\dots,a_{d})$ within partition $k$, the log-density term is expanded as:
$\log\hat{f}_{n,\ell}(\mathbf{X}_{v})=-\log\left(\frac{n_{k}}{n}\cdot\prod_{j=1}^{d}\frac{2m_{k}}{n_{k}\Delta x_{j,a_{j}}^{k}}\right)$
where $\Delta x_{j,a_{j}}^{k}=x_{j,(a_{j}+m_{k})}^{k}-x_{j,(a_{j}-m_{k})}^{k}$. The final estimator is the average of these terms over all $n$ data points.
Building on Theorem 4, the convergence of the entropy estimator can now be established. The proof follows analogously to that of the bivariate entropy consistency. This final theoretical result is stated in the following theorem.
Theorem 5 (Consistency of the Entropy Estimator $\hat{H}_{n,\ell}$).
Assume that the true density $f$ satisfies $\log f\in L^{1+\delta}$ for some $\delta\in(0,1)$, $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$, with $n_{k}\geq 1$, and that the number of partitions satisfies $\ell(n)^{d}=o(n)$. Then, the entropy estimator $\hat{H}_{n,\ell}$ converges in $L^{1}$ to the true entropy $H(f):=\mathbb{E}[-\log f(X)]$ as $n\to\infty$:
$\hat{H}_{n,\ell}\xrightarrow{L^{1}}H(f),\quad\text{and}\quad\hat{H}_{n,\ell}\xrightarrow{\text{a.s.}}H(f).$
Building upon the $L^{1}$ and almost sure convergence of the PSS entropy estimator established in Theorem 5, we can immediately obtain the consistency of the corresponding mutual information estimator.
Corollary 2 (Consistency of the PSS Mutual Information Estimator).
Let $X\in\mathbb{R}^{d_{X}}$ and $Y\in\mathbb{R}^{d_{Y}}$ be random vectors. Define the PSS mutual information estimator as
$\hat{I}_{n,\ell}(X;Y)\;=\;\hat{H}_{n,\ell}(X)+\hat{H}_{n,\ell}(Y)-\hat{H}_{n,\ell}(X,Y).$
Under the same assumptions as Theorem 5,
$\hat{I}_{n,\ell}(X;Y)\xrightarrow{L^{1}}I(X;Y),\ \hat{I}_{n,\ell}(X;Y)\xrightarrow{a.s.}I(X;Y)$
5 Numerical Experiments
We compare PSS against kkNN estimators (KL, KSG), CADEE, and uniformizing-mapping versions (UM-tKL, UM-tKSG). To ensure a fair comparison, the key hyperparameter for each method—the partition parameter $\ell$ for PSS and the number of neighbors $k$ for kNN estimators—was chosen to minimize the empirical Mean Squared Error (MSE) in each setting. All results are averaged over 100 trials and reported as rooted mean squared error (RMSE) relative to the true entropy. The standard kkNN estimator has a naive time complexity of $O(N^{2}d)$ in $d$ dimensions, since all pairwise distances are computed. Tree-based methods (k-d tree, Ball tree) reduce query cost to $O(Nd\log N)$ after an $O(Nd\log N)$ build phase, but their efficiency deteriorates with dimension and skewness. Normalizing flow (NF) estimators AoLi2022 require training a deep generative model with $O(Nd)$ forward/backward passes per epoch, repeated over many epochs, leading to substantial computational overhead. In contrast, PSS assigns $N$ samples to $\ell^{d}$ partitions in $O(Nd)$, sorts within partitions in about $O(dN\log(N/\ell^{d}))$, and computes entropy in $O(Nd)$. As $\ell$ grows with $N$, the factor $N/\ell^{d}$ shrinks and the cost approaches $O(Nd)$, yielding favorable scalability. CADEE copula runs in reasonable time for small problems but its runtime grows rapidly with NN and $d$ relative to PSS, and its accuracy degrades in higher dimensions.
Multivariate Normal.
We first consider multivariate normal data with $\rho=0$. Figure 2 reports RMSE and runtime across varying sample sizes ($d=10$) and dimensions ($N=3000$). PSS achieves consistently lower RMSE than kkNN and CADEE, with the advantage widening as $d$ grows. In terms of runtime, PSS scales nearly linearly with NN and remains substantially faster than CADEE, while also outperforming kkNN methods. These results confirm that PSS mitigates the curse of dimensionality in Gaussian settings while maintaining computational efficiency.
Figure 2: RMSE (a, c) and runtime (b, d) for entropy estimators under the Normal distribution with $\rho=0$. Panels (a, b): $d=10$, varying sample size NN. Panels (c, d): $N=3000$, varying dimension.
Figure 3: RMSE (a, c) and runtime (b, d) for entropy estimators under the Gamma distribution with shape =0.4=0.4, scale =0.3=0.3. Panels (a, b): $d=5$, varying sample size NN. Panels (c, d): $\rho=0$ $N=30000$, varying dimension.
Multivariate Gamma.
We next evaluate Gamma marginals combined through a Gaussian copula with $\rho=0$, shape $0.4$, and scale $0.3$ (Figure 3). In all Gaussian–copula experiments, the oracle joint entropy is computed analytically as
$H(X)\;=\;\sum_{j=1}^{d}h(X_{j})+\tfrac{1}{2}\log\det R,$
where $h(X_{j})$ is the marginal Gamma entropy and $R$ is the copula correlation matrix. PSS remains highly accurate under skewed and heavy-tailed marginals, substantially outperforming kkNN and CADEE at moderate $d$. Runtime comparisons again highlight the scalability of PSS relative to CADEE, whose cost grows steeply with NN and $d$. These findings demonstrate that PSS adapts effectively to skewness and bounded supports while preserving its computational advantage.
6 Discussion
We introduced the partitioned sample-spacing (PSS) estimator, a nonparametric method for joint entropy. Theoretical guarantees include almost sure and $L^{1}$ consistency. Empirically, PSS outperforms kkNN and related methods in moderate to high dimensions and adapts well to skewed distributions, while scaling efficiently with sample size. Its efficiency stems from local partitioning, which avoids exhaustive pairwise distance computations. The partition parameter $\ell$ is central: larger $\ell$ improves accuracy at modest additional cost, though automated selection remains an open problem. As with other grid-based methods, the exponential growth of partitions limits applicability in very high dimensions. Indeed, as seen in Figures 2c and 3c, performance can degrade for larger values of l in very high dimensions $(d>20)$, suggesting that overly fine partitions may lead to unstable local estimates in sparse settings.
A natural avenue for addressing the choice of $\ell$ is cross-validation. One may split the data into training and validation folds, fit $\hat{f}_{n,\ell}$ on the training fold, and evaluate the average negative log-density on the validation fold. Since our entropy estimator itself is defined as an empirical average of $\log\hat{f}_{n,\ell}(X)$, this validation criterion is directly aligned with the estimation objective. In effect, CV minimizes a finite-sample surrogate of the Kullback–Leibler risk, suggesting that the selected $\ell$ may asymptotically approach the optimal rate. Compared to kkNN-based estimators, which often require ad hoc bias–variance heuristics, and NF-based estimators, which demand expensive model training, PSS admits a simple and principled likelihood-based validation strategy. Developing formal guarantees for such data-driven selection remains an important future direction.
Overall, PSS offers a consistent, scalable alternative to kkNN. Future work includes characterizing its bias–variance trade-off and developing more principled strategies for tuning $\ell$, which would enhance its utility in machine learning and information-theoretic applications.
7 Additional Experiment: Evaluating Independent Component Analysis on Real-World EEG Data via Cross-Validation
To demonstrate the robustness and reliability of our proposed Partitioned Sample-Spacing (PSS) estimator on real-world, moderate-dimensional data, we conducted an experiment evaluating the performance of Independent Component Analysis (ICA). The procedure is summarized as follows.
1. Data and Preprocessing.
We used the UCI EEG Eye State dataset, a 14-dimensional time-series signal comprising N=14,980N=14{,}980 samples and $d=14$ channels. The dataset is a continuous recording from a single subject. To prepare the data for entropy estimation—which assumes independent and identically distributed (i.i.d.) samples—we first pre-processed the entire time series via whitening, which zero-centers the data and applies a linear transformation to decorrelate the features. Each row of the transformed data matrix was then treated as an i.i.d. sample from a stationary $d$-dimensional distribution. This simplification ignores temporal dependence and is adopted solely for evaluating entropy estimators on real-world correlated data.
2. Analysis Procedure (ICA).
We applied the FastICA algorithm to the whitened data ($\mathbf{X}_{w}$) to separate it into a set of statistically independent components ($\mathbf{Y}$). All computations were performed using scikit-learn 1.5.2 (FastICA, whiten=True) on an Apple M2, macOS 14.6, Python 3.9 environment.
3. Evaluation Metric (Total Correlation).
To quantify the performance of ICA, we measure the Total Correlation (TC), which represents the total statistical dependence in a random vector $\mathbf{X}=(X_{1},\ldots,X_{d})$. It is defined as the sum of the marginal entropies minus the joint entropy:
$\mathrm{TC}(\mathbf{X})=\sum_{i=1}^{d}H(X_{i})-H(\mathbf{X})\,.$
TC is non-negative, with $\mathrm{TC}=0$ indicating perfect independence. A successful ICA should yield components with a TC value close to zero. We measured the TC of the data before applying ICA ($\texttt{TC}_{\text{before}}$) and after ($\texttt{TC}_{\text{after}}$). An estimator’s reliability is judged by its ability to report a small, positive $\texttt{TC}_{\text{after}}$ value.
4. Hyperparameter Tuning.
For a fair comparison, the key hyperparameter for each estimator was selected via 3-fold cross-validation ($K=3$) on the whitened data. As proposed in the discussion section, we employed a principled, likelihood-based validation strategy for PSS: selecting the number of partitions ($\ell$) that minimized the average negative log-density on the held-out folds. This criterion is directly aligned with the entropy estimation objective, as the estimator itself is defined as an empirical average of the negative log-density. This single, globally-optimized hyperparameter, denoted $\ell^{*}$, was then used for all subsequent PSS-based calculations, including both the joint and the marginal entropy estimations. For KL/KSG, we minimize the held-out negative log-density estimated.
The results are visualized in Figure 4 and summarized in Table 1. These highlight the performance of PSS relative to several widely used $k$-NN based estimators, including KL, KSG, and their truncated variants (tKL, tKSG).
Table 1: Summary of ICA performance evaluation by entropy estimator. All TC values are in nats. $\Delta\mathrm{TC}=\texttt{TC}_{\text{before}}-\texttt{TC}_{\text{after}}$.
[…]
PSS $\ell^{*}=12$ 1.439 6.758 1.11 Reliable and Sensitive
KSG $k^{*}=1$ 8.499 3.241 1.91 Moderate Sensitivity
tKSG $k^{*}=1$ 8.084 3.434 2.99 Moderate Sensitivity
KL $k^{*}=1$ 7.877-7.877 17.881 1.14 Violates $\mathrm{TC}\geq 0$ (Estimation Error)
tKL $k^{*}=1$ 10.697-10.697 18.651 0.34 Violates $\mathrm{TC}\geq 0$ (Estimation Error)
[…]
Figure 4: Total Correlation values before ($\texttt{TC}_{\text{before}}$, hollow) and after ICA ($\texttt{TC}_{\text{after}}$, filled) for each estimator. Negative post-ICA TC values indicate estimator bias or variance, not true independence. PSS is the only estimator reporting a stable, low, and nonnegative $\texttt{TC}_{\text{after}}$, aligning with theoretical expectations.
8 Proofs
8.1 Proof of lemma 1
We first claim $X_{(i+m)}-X_{(i)}\xrightarrow{a.s.}0$. Let $\hat{F}_{n}(x)$ be an empirical cdf of $F(X)$. Then, $\hat{F}_{n}(X_{(i+m)})=\frac{1}{n}\sum_{j=1}^{n}I(X_{j}\leq X_{(i+m)})=\frac{i+m}{n}$. Note that $\sup\limits_{x}|F(x)-\hat{F}_{n}(x)|\xrightarrow{a.s.}0$ by Glivenko-Cantelli theorem. Then, $|F(X_{(i)})-F(X_{(i+m)})|$ is bounded by:
$|F(X_{(i)})-\hat{F}_{n}(X_{(i)})|+|\hat{F}_{n}(X_{(i)})-\hat{F}_{n}(X_{(i+m)})|+|\hat{F}_{n}(X_{(i+m)})-F(X_{(i+m)})|$
The first and third term goes to 0 almost surely by Glivenko-Cantelli theorem. The second term goes to 0 deterministically as $|\hat{F}_{n}(X_{(i)})-\hat{F}_{n}(X_{(i+m)})|=\frac{m}{n}$ which goes to 0 by the assumption. Hence,
$F(X_{(i+m)})-F(X_{(i)})\xrightarrow{a.s.}0$
Since $F$ is continuous, $X_{(i+m)}-X_{(i)}\xrightarrow{a.s.}0$. Similarly, one can show that $X_{(i-m)}-X_{(i)}\xrightarrow{a.s.}0$. Since $\xi_{i}=\frac{1}{2m}\sum\limits_{m^{\prime}=-m}^{m-1}X_{(i+m^{\prime})}$,
$|\xi_{i}-X_{(i)}|\ =|\frac{1}{2m}\sum\limits_{m^{\prime}=-m}^{m-1}(X_{(i+m^{\prime})}-X_{(i)})|$ $\leq\frac{1}{2m}\sum\limits_{m^{\prime}=-m}^{m-1}|X_{(i+m^{\prime})}-X_{(i)}|$
$\leq\frac{1}{2m}\sum\limits_{m^{\prime}=-m}^{m-1}|X_{(i+m)}-X_{(i-m)}|$
$=|X_{(i+m)}-X_{(i-m)}|$
As $X_{(i+m)}-X_{(i-m)}\xrightarrow{a.s.}0$, $\xi_{i}-X_{(i)}\xrightarrow{a.s.}0$. Similarly, as $\xi_{i+1}=\frac{1}{2m}\sum\limits_{m^{\prime}=-m+1}^{m}X_{(i+m^{\prime})}$, $\xi_{i+1}-X_{(i)}\xrightarrow{a.s.}0$.
8.2 Proof of Theorem 1
The estimator for $x\in(\xi_{i},\xi_{i+1}]$ can be written as the product of two terms:
$\hat{f}_{n}(x)=\frac{2m/n}{L_{n}}=\left(\frac{2m/n}{P_{n}}\right)\cdot\left(\frac{P_{n}}{L_{n}}\right),$
where $L_{n}=X_{(i+m)}-X_{(i-m)}$ and $P_{n}=F(X_{(i+m)})-F(X_{(i-m)})$. The proof proceeds by showing that the first term in the product converges to 1 almost surely, while the second term converges to $f(x)$ almost surely.
Convergence of $P_{n}$
Let $U_{k}=F(X_{(k)})$. Since $F$ is continuous, $U_{k}$ are the order statistics from a Uniform(0,1) distribution. Thus, $P_{n}=U_{(i+m)}-U_{(i-m)}$. The random variable $P_{n}=U_{(i+m)}-U_{(i-m)}$ follows a Beta distribution, specifically $P_{n}\sim\mathrm{Beta}(2m,n-2m+1)$ david2004order.
This is a standard result for differences of uniform order statistics. If $U_{(1)}<\dots<U_{(n)}$ are order statistics from $U(0,1)$, then for $j>k$, $U_{(j)}-U_{(k)}\sim\mathrm{Beta}(j-k,n-(j-k)+1)$. In our case, we have $U_{(i+m)}-U_{(i-m)}$. Let $j^{\prime}=i+m$ and $k^{\prime}=i-m$. The first parameter is $\alpha=j^{\prime}-k^{\prime}=(i+m)-(i-m)=2m$. The second parameter is $\beta=n-(j^{\prime}-k^{\prime})+1=n-2m+1$. Thus, $P_{n}\sim\mathrm{Beta}(2m,n-2m+1)$. Let $\mu_{n}$ be the mean of $P_{n}$, $\mathbb{E}[P_{n}]=\frac{2m}{2m+(n-2m+1)}=\frac{2m}{n+1}$.
We will now prove that the ratio $\frac{P_{n}}{\mu_{n}}$ converges to 1 almost surely. To do so, we will use the Borel-Cantelli Lemma, which requires us to show that the sum of the probabilities of deviation, $\sum_{n=1}^{\infty}\mathbb{P}\left(\left|\frac{P_{n}}{\mu_{n}}-1\right|>\epsilon\right)$, is finite for any arbitrary $\epsilon>0$. We employ a standard concentration inequality for the Beta distribution, which is sharper than the general Hoeffding bound. For a random variable $X\sim\text{Beta}(\alpha,\beta)$ with mean $\mu$, a bound can be derived from the Chernoff method Massart. One such bound is given by:
$\mathbb{P}(|X-\mu|\geq t)\leq 2\exp\left(-\frac{(\alpha+\beta)t^{2}}{2\mu(1-\mu)+\frac{2}{3}t}\right)$
For our problem, we apply this with $X=P_{n}$, $\alpha+\beta=n+1$, $\mu=\mu_{n}$, and $t=t_{n}=\mu_{n}\epsilon$. The probability of the event $A_{n}(\epsilon)=\{|P_{n}-\mu_{n}|>\mu_{n}\epsilon\}$ is thus bounded by:
$\mathbb{P}(A_{n}(\epsilon))\leq 2\exp\left(-\frac{(n+1)(\mu_{n}\epsilon)^{2}}{2\mu_{n}(1-\mu_{n})+\frac{2}{3}\mu_{n}\epsilon}\right)$ $=2\exp\left(-\frac{2m\epsilon^{2}}{2(1-\mu_{n})+\frac{2}{3}\epsilon}\right)$
$\leq 2\exp\left(-\frac{\epsilon^{2}\sqrt{n}}{1+\frac{1}{3}\epsilon}\right)$
Let $K=\frac{\epsilon^{2}}{1+1/3\epsilon}$. We use the Limit Comparison Test with the convergent p-series $\sum_{n=1}^{\infty}\frac{1}{n^{2}}$. We evaluate the limit of the ratio:
$\lim_{n\to\infty}\frac{2\exp(-K\sqrt{n})}{1/n^{2}}=2\lim_{n\to\infty}\frac{n^{2}}{e^{K\sqrt{n}}}=0$
Since the limit is 0 and $\sum\frac{1}{n^{2}}$ converges, the series $\sum_{n=1}^{\infty}2\exp(-K\sqrt{n})$ also converges.
Hence,
$\sum_{n=1}^{\infty}\mathbb{P}\left(\left|\frac{P_{n}}{\mu_{n}}-1\right|>\epsilon\right)=$ $\sum_{n=1}^{\infty}\mathbb{P}\left(\left|P_{n}-\mu_{n}\right|>\mu_{n}\epsilon\right)$
$=$ $\sum_{n=1}^{\infty}\mathbb{P}(A_{n}(\epsilon))\leq\sum_{n=1}^{\infty}2\exp(-K\sqrt{n})<\infty$
By Borel-Cantelli Lemma, $\frac{P_{n}}{\mu_{n}}\xrightarrow{a.s.}1$. Its reciprocal $\frac{\mu_{n}}{P_{n}}$ also converges to 1 almost surely by the continuous mapping theorem. Finally, since the deterministic sequence $\frac{n+1}{n}$ converges to 1, the almost sure convergence of $\frac{\mu_{n}}{P_{n}}$ to 1 directly implies the convergence of their product:
$\frac{2m/n}{P_{n}}=\left(\frac{\mu_{n}}{P_{n}}\right)\cdot\left(\frac{n+1}{n}\right)\xrightarrow{a.s.}1$
Convergence of $\frac{P_{n}}{L_{n}}$ to $f(x)$
We have $\frac{P_{n}}{L_{n}}=\frac{F(X_{(i+m)})-F(X_{(i-m)})}{X_{(i+m)}-X_{(i-m)}}$. Then by the Mean Value Theorem, there exists $X^{\prime}_{i}\in(X_{(i-m)},X_{(i+m)})$ such that
$\frac{F(X_{(i+m)})-F(X_{(i-m)})}{X_{(i+m)}-X_{(i-m)}}=f(X^{\prime}_{i})$
Since $x\in(\xi_{i},\xi_{i+1}]$, and $\xi_{i+1}-\xi_{i}\xrightarrow{a.s.}0$ (by Lemma 1), for $x$ to remain in this asymptotically vanishing interval, it must be that $\xi_{i}\xrightarrow{a.s.}x$. Given $\xi_{i}\xrightarrow{a.s.}X_{(i)}$ and $\xi_{i}\xrightarrow{a.s.}x$, it follows that $X_{(i)}\xrightarrow{a.s.}x$.
Now, using the results from Lemma 1 regarding $X_{(i-m)}$ and $X_{(i+m)}$:
Since $X_{(i-m)}\xrightarrow{a.s.}X_{(i)}$ and $X_{(i)}\xrightarrow{a.s.}x$, by transitivity of almost sure convergence, $X_{(i-m)}\xrightarrow{a.s.}x$. Similarly, since $X_{(i+m)}\xrightarrow{a.s.}X_{(i)}$ and $X_{(i)}\xrightarrow{a.s.}x$, then $X_{(i+m)}\xrightarrow{a.s.}x$.
We have $X_{(i-m)}\leq X^{\prime}_{i}\leq X_{(i+m)}$. Since $X_{(i-m)}\xrightarrow{a.s.}x$ and $X_{(i+m)}\xrightarrow{a.s.}x$, by the squeeze theorem, $X^{\prime}_{i}\xrightarrow{a.s.}x$. Since $f$ is continuous at $x$, by the Continuous Mapping Theorem, $f(X^{\prime}_{i})\xrightarrow{a.s.}f(x)$. Therefore, $\frac{P_{n}}{L_{n}}\xrightarrow{a.s.}f(x)$.
We have $\hat{f}_{n}(x)=\left(\frac{2m/n}{P_{n}}\right)\cdot\left(\frac{P_{n}}{L_{n}}\right)$. Since $\frac{2m/n}{P_{n}}\xrightarrow{a.s.}1$ and $\frac{P_{n}}{L_{n}}\xrightarrow{a.s.}f(x)$, their product converges almost surely:
$\hat{f}_{n}(x)\xrightarrow{a.s.}1\cdot f(x)=f(x)$
8.3 Proof of Proposition 1
Let the sample space be divided into $\ell^{2}$ partitions. The partitioned density estimator $\hat{f}_{n,\ell}$ in a cell $(\xi_{i}^{k},\xi_{i+1}^{k}]\times(\eta_{j}^{k},\eta_{j+1}^{k}]$ within partition $P_{k}$ is defined as
$\hat{f}_{n,\ell}(x,y)=\frac{n_{k}}{n}\cdot\frac{4m_{k}^{2}}{n_{k}^{2}\Delta x_{i}^{k}\Delta y_{j}^{k}}$
where $\Delta x_{i}^{k}=x_{(i+m_{k})}^{k}-x_{(i-m_{k})}^{k}$ and $\Delta y_{j}^{k}=y_{(j+m_{k})}^{k}-y_{(j-m_{k})}^{k}$. The total integral is the sum of integrals over each partition. The integral over a single partition $P_{k}$ can be split into contributions from its $(n_{k}-1)^{2}$ interior cells and its boundary cells: $\int_{C_{k}}\hat{f}_{n,\ell}=InteriorMass_{k}+EdgeMass_{k}$. For any interior cell, the ratio of the grid spacing to the m-spacing simplifies. Using the definition $\eta_{j}^{k}=\frac{1}{2m_{k}}\sum_{l=j-m_{k}}^{j+m_{k}-1}y_{(l)}^{k}$, the difference $\eta_{j+1}^{k}-\eta_{j}^{k}$ becomes a telescoping sum:
$\eta_{j+1}^{k}-\eta_{j}^{k}$ $=\frac{1}{2m_{k}}\left(\sum_{l=j-m_{k}+1}^{j+m_{k}}y_{(l)}^{k}-\sum_{l=j-m_{k}}^{j+m_{k}-1}y_{(l)}^{k}\right)=\frac{1}{2m_{k}}\left(y_{(j+m_{k})}^{k}-y_{(j-m_{k})}^{k}\right)=\frac{\Delta y_{j}^{k}}{2m_{k}}$
Thus, the ratio $\frac{\eta_{j+1}^{k}-\eta_{j}^{k}}{\Delta y_{j}^{k}}$ is exactly $\frac{1}{2m_{k}}$. Similarly, $\frac{\xi_{i+1}^{k}-\xi_{i}^{k}}{\Delta x_{i}^{k}}=\frac{1}{2m_{k}}$. Then, The interior mass is:
$InteriorMass_{k}=$ $\sum_{i=1}^{n_{k}-1}\sum_{j=1}^{n_{k}-1}\left(\frac{n_{k}}{n}\frac{4m_{k}^{2}}{n_{k}^{2}\Delta x_{i}^{k}\Delta y_{j}^{k}}\right)\times(\xi_{i+1}^{k}-\xi_{i}^{k})(\eta_{j+1}^{k}-\eta_{j}^{k})$
$=$ $\frac{n_{k}}{n}\frac{4m_{k}^{2}}{n_{k}^{2}}\left(\sum_{i=1}^{n_{k}-1}\frac{\xi_{i+1}^{k}-\xi_{i}^{k}}{\Delta x_{i}^{k}}\right)\left(\sum_{j=1}^{n_{k}-1}\frac{\eta_{j+1}^{k}-\eta_{j}^{k}}{\Delta y_{j}^{k}}\right)=\frac{n_{k}}{n}\left(1-\frac{1}{n_{k}}\right)^{2}$
The edge mass depends on boundary terms $E_{x}^{k}:=\frac{\xi_{1}^{k}-\xi_{0}^{k}}{\Delta x_{0}^{k}}+\frac{\xi_{n_{k}+1}^{k}-\xi_{n_{k}}^{k}}{\Delta x_{n_{k}}^{k}}$ and $E_{y}^{k}:=\frac{\eta_{1}^{k}-\eta_{0}^{k}}{\Delta y_{0}^{k}}+\frac{\eta_{n_{k}+1}^{k}-\eta_{n_{k}}^{k}}{\Delta y_{n_{k}}^{k}}$. Using the definitions of $\xi_{i}^{k}$ and the boundary conditions on order statistics, we can show $0\leq E_{x}^{k},E_{y}^{k}\leq 1$. The first term of $E_{x}^{k}$ is bounded as:
$\frac{\xi_{1}^{k}-x_{(1)}^{k}}{\Delta x_{0}^{k}}=\frac{\sum_{j=1}^{m_{k}}(x_{(j)}^{k}-x_{(1)}^{k})}{2m_{k}\Delta x_{0}^{k}}\leq\frac{m_{k}(x_{(m_{k})}^{k}-x_{(1)}^{k})}{2m_{k}\Delta x_{0}^{k}}=\frac{1}{2}$
A similar argument for the second term shows $E_{y}^{k}\leq 1/2+1/2=1$. The edge mass denoted as $EdgeMass_{k}$ is the sum of the remaining terms from the factorization:
$\frac{n_{k}}{n}\frac{4m_{k}^{2}}{n_{k}^{2}}\left(\sum_{i=1}^{n_{k}-1}\frac{\xi_{i+1}^{k}-\xi_{i}^{k}}{\Delta x_{i}^{k}}\right)E_{y}^{k}+\frac{n_{k}}{n}\frac{4m_{k}^{2}}{n_{k}^{2}}\left(\sum_{j=1}^{n_{k}-1}\frac{\eta_{j+1}^{k}-\eta_{j}^{k}}{\Delta y_{j}^{k}}\right)E_{x}^{k}+\frac{n_{k}}{n}\frac{4m_{k}^{2}}{n_{k}^{2}}E_{x}^{k}E_{y}^{k}$
$=$ $\frac{n_{k}}{n}\frac{4m_{k}^{2}}{n_{k}^{2}}\left[\left(\frac{n_{k}-1}{2m_{k}}\right)E_{y}^{k}+\left(\frac{n_{k}-1}{2m_{k}}\right)E_{x}^{k}+E_{x}^{k}E_{y}^{k}\right]$
$\leq$ $\frac{n_{k}}{n}\left[\frac{2m_{k}(n_{k}-1)}{n_{k}^{2}}+\frac{2m_{k}(n_{k}-1)}{n_{k}^{2}}+\frac{4m_{k}^{2}}{n_{k}^{2}}\right]=\frac{n_{k}}{n}O\left(\frac{m_{k}}{n_{k}}\right)$
Summing over all partitions:
$\int_{\mathbb{R}^{2}}\hat{f}_{n,\ell}$ $=\sum_{k=1}^{\ell^{2}}InteriorMass_{k}+\sum_{k=1}^{\ell^{2}}EdgeMass_{k}$
$=\sum_{k=1}^{\ell^{2}}\frac{n_{k}}{n}\left(1-\frac{1}{n_{k}}\right)^{2}+\sum_{k=1}^{\ell^{2}}\frac{n_{k}}{n}O\left(\frac{m_{k}}{n_{k}}\right)$
$=1-O\left(\frac{\ell^{2}}{n}\right)+O\left(\max_{k}\frac{m_{k}}{n_{k}}\right)$
As $n\to\infty$, since $\ell(n)^{2}=o(n)$ and $m_{k}/n_{k}\to 0$, the total integral converges to 1.
8.4 Proof of lemma 2
We aim to show that $\hat{f}_{X}^{k}(x)\xrightarrow{a.s.}f_{X|C_{k}}(x)$ for $x\in(\xi_{i}^{k},\xi_{i+1}^{k}]$, where $f_{X|C_{k}}(x)$ is the true conditional probability density function of $X$ given partition $P_{k}$. The estimator is
$\hat{f}_{X}^{k}(x)=\left(\frac{2m_{k}/n_{k}}{P_{n}^{k}}\right)\cdot\left(\frac{P_{n}^{k}}{L_{n}^{k}}\right)$
where $P_{n}^{k}=F_{X|C_{k}}(X_{(i+m_{k})}^{k})-F_{X|C_{k}}(X_{(i-m_{k})}^{k})$ and $L_{n}^{k}=X_{(i+m_{k})}^{k}-X_{(i-m_{k})}^{k}$. The samples $X_{(j)}^{k}$ are the $n_{k}$ order statistics from partition $P_{k}$. This proof relies on the conditions $n_{k}\to\infty$, $m_{k}\to\infty$, $m_{k}/n_{k}\to 0$ as $n\to\infty$, and continuity of $f_{X|C_{k}}(x)$ at $x$.
The argument closely follows that of Theorem 1.
Convergence of the first term: Let $U_{(j)}^{k}=F_{X|C_{k}}(X_{(j)}^{k})$. These are the order statistics from a Uniform(0,1) distribution, based on the $n_{k}$ samples in $C_{k}$. Thus, $P_{n}^{k}=U_{(i+m_{k})}^{k}-U_{(i-m_{k})}^{k}\sim\mathrm{Beta}(2m_{k},n_{k}-2m_{k}+1)$. Let $\mu_{n}^{k}$ be expectation $\mathbb{E}[P_{n}^{k}]=\frac{2m_{k}}{n_{k}+1}$. Then,
$\sum_{n=1}^{\infty}\mathbb{P}\left(\left|\frac{P_{n}^{k}}{\mu_{n}^{k}}-1\right|>\epsilon\right)=\sum_{n=1}^{\infty}\mathbb{P}\left(\left|P_{n}^{k}-\mu_{n}^{k}\right|>\mu_{n}^{k}\epsilon\right)\leq\sum_{n=1}^{\infty}2\exp\left(-\frac{\epsilon^{2}\sqrt{n_{k}}}{1+\frac{1}{3}\epsilon}\right)<\infty$
By Borel-Cantelli Lemma, $\frac{P_{n}^{k}}{\mu_{n}^{k}}\xrightarrow{a.s.}1$. Its reciprocal $\frac{\mu_{n}^{k}}{P_{n}^{k}}$ also converges to 1 almost surely by the continuous mapping theorem. Finally, since the deterministic sequence $\frac{n_{k}+1}{n_{k}}$ converges to 1, $\frac{2m_{k}/n_{k}}{P_{n}^{k}}\xrightarrow{a.s.}1$
Convergence of the second term: By the Mean Value Theorem (assuming $F_{X|C_{k}}$ is differentiable with derivative $f_{X|C_{k}}$):
$\frac{P_{n}^{k}}{L_{n}^{k}}=\frac{F_{X|C_{k}}(X_{(i+m_{k})}^{k})-F_{X|C_{k}}(X_{(i-m_{k})}^{k})}{X_{(i+m_{k})}^{k}-X_{(i-m_{k})}^{k}}=f_{X|C_{k}}(X^{\prime}_{k})$
for some $X^{\prime}_{k}\in(X_{(i-m_{k})}^{k},X_{(i+m_{k})}^{k})$. We need to show $X^{\prime}_{k}\xrightarrow{a.s.}x$. The interval $(\xi_{i}^{k},\xi_{i+1}^{k}]$, in which $x$ lies, is defined using averages of order statistics $X_{(j)}^{k}$ partition $k$, similar to $\xi_{i}$ in Lemma 1. Applying the logic of Lemma 1 to the $n_{k}$ samples within partition $k$: $X_{(i\pm m_{k})}^{k}\xrightarrow{a.s.}X_{(i)}^{k}$ (since $m_{k}/n_{k}\to 0$). Furthermore, for $x\in(\xi_{i}^{k},\xi_{i+1}^{k}]$, $X_{(i)}^{k}$ is the relevant order statistic that converges almost surely to $x$. Since $X^{\prime}_{k}$ is between $X_{(i-m_{k})}^{k}$ and $X_{(i+m_{k})}^{k}$, both of which converge almost surely to $x$, it follows by the Sandwich Theorem that $X^{\prime}_{k}\xrightarrow{a.s.}x$. Given that $f_{X|C_{k}}$ is continuous at $x$, $f_{X|C_{k}}(X^{\prime}_{k})\xrightarrow{a.s.}f_{X|C_{k}}(x)$ by the Continuous Mapping Theorem. Finally, the product of the two terms converges almost surely:
$\hat{f}_{X}^{k}(x)=\left(\frac{2m_{k}/n_{k}}{P_{n}^{k}}\right)\cdot\left(\frac{P_{n}^{k}}{L_{n}^{k}}\right)\xrightarrow{a.s.}1\cdot f_{X|C_{k}}(x)=f_{X|C_{k}}(x)$
8.5 Proof of Theorem 2
Let the partition $k$ containing the point $(x,y)$ be defined by the rectangular region $x_{\text{min},k},x_{\text{max},k}]\times[y_{\text{min},k},y_{\text{max},k}]$. Let $\Delta x_{k}=x_{\text{max},k}-x_{\text{min},k}$ be the width and $\Delta y_{k}=y_{\text{max},k}-y_{\text{min},k}$ be the height of this partition, so its area is $A_{k}=\Delta x_{k}\Delta y_{k}$. The estimator for $f(x,y)$ at the point $(x,y)$ is: $\hat{f}_{n,l}(x,y)=\frac{n_{k}}{n}\cdot\hat{f}_{X}^{k}(x)\cdot\hat{f}_{Y}^{k}(y)$. We aim to show that $\hat{f}_{n,l}(x,y)\xrightarrow{a.s.}f(x,y)$ as $n,\ell\to\infty$.
Convergence of the Sample Proportion $\frac{n_{k}}{n}$
Let $P(C_{k})=\iint_{k}f(u,s)\,du\,ds=\int_{y_{\text{min},k}}^{y_{\text{max},k}}\int_{x_{\text{min},k}}^{x_{\text{max},k}}f(u,s)\,du\,ds$ be the true probability mass of partition $k$. Because the indicator of falling into the partition $k$ is Bernoulli with mean $P(C_{k})$ by Kolmogorov’s Strong Law of Large Numbers, the sample proportion $\frac{n_{k}}{n}$ converges almost surely to $P(C_{k})$. This relies on the conditions in Lemma 2 that ensure $n_{k}\to\infty$ for the shrinking partition $k$ containing $(x,y)$.
Consistency of Univariate Density Estimators $\hat{f}_{X}^{k}(x)$ and $\hat{f}_{Y}^{k}(y)$
The true conditional PDF of $X$ at $x$, given $C_{k}$, is derived as follows: The conditional PDF $f_{X|C_{k}}(x)$ is defined from the conditional CDF $F_{X|C_{k}}(x)=P(X\leq x|C_{k})$ as:
$f_{X|C_{k}}(x)=\frac{d}{dx}P(X\leq x|C_{k})=\frac{1}{P(C_{k})}\frac{d}{dx}P(X\leq x,(X,Y)\in C_{k})$
The joint probability $P(X\leq x,(X,Y)\in C_{k})$ is:
$P(X\leq x,(X,Y)\in C_{k})=\int_{y_{\text{min},k}}^{y_{\text{max},k}}\left(\int_{x_{\text{min},k}}^{x}f(u,s)\,du\right)ds$
for $x_{\text{min},k}\leq x\leq x_{\text{max},k}$. Differentiating with respect to $x$:
$\frac{d}{dx}P(X\leq x,(X,Y)\in C_{k})$ $=\int_{y_{\text{min},k}}^{y_{\text{max},k}}\left(\frac{d}{dx}\int_{x_{\text{min},k}}^{x}f(u,s)\,du\right)ds$
$=\int_{y_{\text{min},k}}^{y_{\text{max},k}}f(x,s)\,ds$
by the Fundamental Theorem of Calculus. Thus,
$f_{X|C_{k}}(x)=\frac{\int_{y_{\text{min},k}}^{y_{\text{max},k}}f(x,s)\,ds}{P(C_{k})}$
Similarly, for $Y$:
$f_{Y|C_{k}}(y)=\frac{\int_{x_{\text{min},k}}^{x_{\text{max},k}}f(u,y)\,du}{P(C_{k})}$
The existence and continuity of these conditional PDFs are ensured by the assumption that the joint PDF $f(x,y)$ is continuous and positive, and $P(C_{k})>0$.
According to Lemma 2, under the conditions on $l(n)$, $n_{k}$, and $m_{k}$:
$\hat{f}_{X}^{k}(x)\xrightarrow{a.s.}f_{X|C_{k}}(x)\qquad\hat{f}_{Y}^{k}(y)\xrightarrow{a.s.}f_{Y|C_{k}}(y)$
Since $\frac{n_{k}}{n}\xrightarrow{a.s.}P(C_{k})$, $\hat{f}_{X;k}(x)\xrightarrow{a.s.}f_{X|C_{k}}(x)$, and $\hat{f}_{Y;k}(y)\xrightarrow{a.s.}f_{Y|C_{k}}(y)$, their product also converges almost surely:
$\hat{f}_{n,l}(x,y)\xrightarrow{a.s.}P(C_{k})\cdot f_{X|C_{k}}(x)\cdot f_{Y|C_{k}}(y)$
Limit of the Target Expression $P(C_{k})\cdot f_{X|C_{k}}(x)\cdot f_{Y|C_{k}}(y)$
Let $Q_{k}(x,y)=P(C_{k})\cdot f_{X|C_{k}}(x)\cdot f_{Y|C_{k}}(y)$. Substituting the integral expressions for the conditional densities:
$Q_{k}(x,y)$ $=P(C_{k})\cdot\left(\frac{\int_{y_{\text{min},k}}^{y_{\text{max},k}}f(x,s)\,ds}{P(C_{k})}\right)\cdot\left(\frac{\int_{x_{\text{min},k}}^{x_{\text{max},k}}f(u,y)\,du}{P(C_{k})}\right)$
$=\frac{1}{P(C_{k})}\left(\int_{y_{\text{min},k}}^{y_{\text{max},k}}f(x,s)\,ds\right)\left(\int_{x_{\text{min},k}}^{x_{\text{max},k}}f(u,y)\,du\right)$
As $n\to\infty$, the conditions (including $l(n)\to\infty$ from Lemma 2) ensure that the partition $k$ containing $(x,y)$ shrinks, so its lengths $\Delta x_{k}\to 0$ and $\Delta y_{k}\to 0$. Since $f(u,s)$ is continuous at $(x,y)$:
By the Mean Value Theorem for Integrals, There exists $\tilde{y}_{k}\in[y_{\text{min},k},y_{\text{max},k}]$ such that $\int_{y_{\text{min},k}}^{y_{\text{max},k}}f(x,s)\,ds=f(x,\tilde{y}_{k})\Delta y_{k}$. Also, there exists $\tilde{x}_{k}\in[x_{\text{min},k},x_{\text{max},k}]$ such that $\int_{x_{\text{min},k}}^{x_{\text{max},k}}f(u,y)\,du=f(\tilde{x}_{k},y)\Delta x_{k}$. Lastly, there exists $(\bar{x}_{k},\bar{y}_{k})\in[x_{\text{min},k},x_{\text{max},k}]\times[y_{\text{min},k},y_{\text{max},k}]$ such that $P(C_{k})=\iint_{k}f(u,s)\,du\,ds=f(\bar{x}_{k},\bar{y}_{k})A_{k}=f(\bar{x}_{k},\bar{y}_{k})\Delta x_{k}\Delta y_{k}$. Substituting these into the expression for $Q_{k}(x,y)$:
$Q_{k}(x,y)=\frac{(f(x,\tilde{y}_{k})\Delta y_{k})(f(\tilde{x}_{k},y)\Delta x_{k})}{f(\bar{x}_{k},\bar{y}_{k})\Delta x_{k}\Delta y_{k}}=\frac{f(x,\tilde{y}_{k})f(\tilde{x}_{k},y)}{f(\bar{x}_{k},\bar{y}_{k})}$
(assuming $f(\bar{x}_{k},\bar{y}_{k})\neq 0$, which is consistent with $f(x,y)>0$ and the shrinking partition).
Now, we take the limit as $n\to\infty$, which implies $\Delta x_{k}\to 0$ and $\Delta y_{k}\to 0$. Since partition $k$ shrinks to the point $(x,y)$: $\tilde{y}_{k}\to y,\tilde{x}_{k}\to x,\bar{x}_{k}\to x\;\text{and}\;\bar{y}_{k}\to y$.
Given that $f(u,s)$ is continuous at $(x,y)$:
$\lim_{n\to\infty}f(x,\tilde{y}_{k})=f(x,y),\ \lim_{n\to\infty}f(\tilde{x}_{k},y)=f(x,y)\ \text{and}\ \lim_{n\to\infty}f(\bar{x}_{k},\bar{y}_{k})=f(x,y)$
Therefore, since $f(x,y)>0$:
$\lim_{n\to\infty}Q_{k}(x,y)=\frac{f(x,y)\cdot f(x,y)}{f(x,y)}=f(x,y)$
We have established that $\hat{f}_{n,l}(x,y)\xrightarrow{a.s.}Q_{k}(x,y)$ and that $Q_{k}(x,y)\to f(x,y)$ as $n\to\infty$ (since $k$ depends on $n$ via $l(n)$ and shrinks to $(x,y)$). Therefore, it follows that:
$\hat{f}_{n,l}(x,y)\xrightarrow{a.s.}f(x,y)$
8.6 Proof of Proposition 2
Let $x_{v}^{\prime}=a\,x_{v}+b,\ y_{v}^{\prime}=c\,y_{v}+d$ for some constant $a,b>0$ and $c,d\in\mathbb{R}$ and note that the grid lines transform as $\eta_{i}^{k^{\prime}}=a\,\eta_{i}^{k}+b$, $\xi_{j}^{k^{\prime}}=c\,\xi_{j}^{k}+d$. Then,
$H_{v}(x^{\prime},y^{\prime},n,\ell)$
$=$ $\frac{1}{n}\sum_{v=1}^{n}\sum_{k=1}^{\ell^{2}}\sum_{j=1}^{n_{k}}\sum_{i=1}^{n_{k}}\log\!\Bigl(\tfrac{n\,n_{k}\,(x^{\prime}_{(i+m_{k})}-x^{\prime}_{(i-m_{k})})\,(y^{\prime}_{(j+m_{k})}-y^{\prime}_{(j-m_{k})})}{4m_{k}^{2}}\Bigr)\,I(\eta_{i}^{k^{\prime}}<x^{\prime}_{v}\leq\eta_{i+1}^{k^{\prime}})\,I(\xi_{j}^{k^{\prime}}<y^{\prime}_{v}\leq\xi_{j+1}^{k^{\prime}})$
$=$ $\frac{1}{n}\sum_{k,v,j,i}\log\!\Bigl(\tfrac{n\,n_{k}\,\bigl(a(x_{(i+m_{k})}-x_{(i-m_{k})})\bigr)\,\bigl(c(y_{(j+m_{k})}-y_{(j-m_{k})})\bigr)}{4m_{k}^{2}}\Bigr)\,I(a\eta_{i}^{k}<ax_{v}\leq a\eta_{i+1}^{k})\,I(c\xi_{j}^{k}<cy_{v}\leq c\xi_{j+1}^{k})$
$=$ $\frac{1}{n}\sum_{k,v,j,i}\Bigl[\log(a\,c)+\log\!\bigl(\tfrac{n\,n_{k}\,(x_{(i+m_{k})}-x_{(i-m_{k})})\,(y_{(j+m_{k})}-y_{(j-m_{k})})}{4m_{k}^{2}}\bigr)\Bigr]\,I(\eta_{i}^{k}<x_{v}\leq\eta_{i+1}^{k})\,I(\xi_{j}^{k}<y_{v}\leq\xi_{j+1}^{k})$
$=$ $\log(a\,c)\;+\;\frac{1}{n}\sum_{k,v,j,i}\log\!\bigl(\tfrac{n\,n_{k}\,(x_{(i+m_{k})}-x_{(i-m_{k})})\,(y_{(j+m_{k})}-y_{(j-m_{k})})}{4m_{k}^{2}}\bigr)\,I(\eta_{i}^{k}<x_{v}\leq\eta_{i+1}^{k})\,I(\xi_{j}^{k}<y_{v}\leq\xi_{j+1}^{k})$
$=$ $\log(a\,c)\;+\;H_{v}(x,y,n,\ell).$
Since $\log(a\,c)$ is independent of $\ell$, it does not affect the location of the minimum. Hence
$\arg\min_{\ell}H_{v}(x^{\prime},y^{\prime},n,\ell)=\arg\min_{\ell}H_{v}(x,y,n,\ell),$
showing that the optimal $\ell$ is invariant under arbitrary shifts and positive scalings of the data. Because any change of mean and variance is just a positive scaling followed by a shift, and under any such affine transform $H$ only gains an additive constant, the $\ell$ that minimizes $H$ must be unchanged. Hence, the optimal partitioning parameter $\ell$ is invariant under mean and variance.
8.7 Proof of Theorem 3
The proof relies on Vitali’s Convergence Theorem via the de la Vallée–Poussin criterion. We have already established from Theorem 2 that $\log\hat{f}_{n,\ell}(X,Y)\xrightarrow{a.s.}\log f(X,Y)$. The remaining critical step is to prove that the sequence of random variables $\{\log\hat{f}_{n,\ell}\}$ is uniformly integrable. We achieve this by showing it is bounded in $L^{1+\delta}$ for some $\delta\in(0,1]$.
The log-density estimator can be decomposed as:
$\log\hat{f}_{n,\ell}(x,y)=\log\left(\frac{n_{k}}{n}\right)+\log\hat{f}_{X}^{k}(x)+\log\hat{f}_{Y}^{k}(y)$
where for a point $(x,y)$ in the sub-grid cell $(\xi_{i}^{k},\xi_{i+1}^{k}]\times(\eta_{j}^{k},\eta_{j+1}^{k}]$, we define $\hat{f}_{X}^{k}(x)=S_{x}T_{x}$ and $\hat{f}_{Y}^{k}(y)=S_{y}T_{y}$. The terms are:
$S_{x}=\frac{2m_{k}}{n_{k}\{F_{X|C_{k}}(x_{(i+m_{k})}^{k})-F_{X|C_{k}}(x_{(i-m_{k})}^{k})\}},\quad T_{x}=\frac{F_{X|C_{k}}(x_{(i+m_{k})}^{k})-F_{X|C_{k}}(x_{(i-m_{k})}^{k})}{x_{(i+m_{k})}^{k}-x_{(i-m_{k})}^{k}}$
and analogously for $S_{y}$ and $T_{y}$. Then, $\log\hat{f}_{n,\ell}=\log(\frac{n_{k}}{n})+\log S_{x}+\log T_{x}+\log S_{y}+\log T_{y}$. By the mean value theorem inside marginal partition interval $I_{y}^{(k)}$ and $I_{x}^{(k)}$ we have, $T_{x}=\frac{|I_{y}^{(k)}|}{p_{k}}\,f(X^{\prime},y_{k}^{*}),T_{y}=\frac{|I_{x}^{(k)}|}{p_{k}}\,f(x_{k}^{*},Y^{\prime})$ for some $X^{\prime},Y^{\prime},y_{k}^{*},x_{k}^{*}\in C_{k}$ where $p_{k}=\mathbb{P}\{X\in C_{k}\}$. Using $\log(n_{k}/(np_{k}))+\log p_{k}=\log(n_{k}/n)$ and $|C_{k}|=|I_{x}^{(k)}||I_{y}^{(k)}|$ we obtain,
$\log\hat{f}_{n,\ell}(X,Y)=\log\frac{n_{k}}{np_{k}}+\log S_{x}+\log S_{y}+\log f(X^{\prime},y_{k}^{*})+\log f(x_{k}^{*},Y^{\prime})+\log\frac{|C_{k}|}{p_{k}}$
By Hölder’s inequality, for any real numbers $x_{1},\dots,x_{m}$ and any $\varepsilon>0$ we have $\left|\sum_{j=1}^{m}x_{j}\right|^{1+\varepsilon}\leq m^{\varepsilon}\sum_{j=1}^{m}|x_{j}|^{1+\varepsilon}$. Thus,
$\mathbb{E}[|\log\hat{f}_{n,\ell}(x,y)|^{1+\delta}]$ $\leq 6^{\delta}\big(\mathbb{E}[|\log\tfrac{n_{k}}{np_{k}}|^{1+\delta}]+\mathbb{E}[|\log S_{x}|^{1+\delta}]+\mathbb{E}[|\log f(X^{\prime},y_{k}^{*})|^{1+\delta}]$
$\quad+\mathbb{E}[|\log S_{y}|^{1+\delta}]+\mathbb{E}[|\log f(x_{k}^{*},Y^{\prime})|^{1+\delta}]+\mathbb{E}[|\log\frac{|C_{k}|}{p_{k}}|^{1+\delta}]\big)$
The proof reduces to showing that each expectation term is uniformly bounded in $n$.
Finiteness of $\mathbb{E}[|\log(n_{k}/np_{k})|^{1+\delta}]$
Let $U_{k}:=\frac{n_{k}-np_{k}}{np_{k}}$ so that $\frac{n_{k}}{np_{k}}=1+U_{k}$. Also, set an event $\mathcal{A}:=\{|U_{k}|\leq\frac{1}{2}\}$. Then, $\mathbb{P}(\mathcal{A}^{c})\leq 2\exp\{-cnp_{k}\}$ by Chernoff bound. Now, by Lyapunov Inequality,
$\mathbb{E}\big[|U_{k}|^{1+\delta}\mid C_{k}\big]\ \leq\ \big(\mathbb{E}[U_{k}^{2}\mid C_{k}]\big)^{(1+\delta)/2}=\Big(\frac{p_{k}(1-p_{k})}{np_{k}^{2}}\Big)^{\!(1+\delta)/2}\ \leq\ C(np_{k})^{-(1+\delta)/2}.$
Similarly $\mathbb{E}[|U_{k}|^{2(1+\delta)}\mid C_{k}]\leq C(np_{k})^{-(1+\delta)}$. Using the inequality for $|u|\leq 1/2$, $|\log(1+u)|\leq C(|u|+u^{2})$, we can bound $\mathbb{E}[|\log{\frac{n_{k}}{n}}|^{1+\delta}]$ as below:
$\mathbb{E}[|\log{\frac{n_{k}}{np_{k}}}|^{1+\delta}]$ $\leq\mathbb{E}[|\log{(1+U_{k})}|^{1+\delta}\mid\mathcal{A}]+\mathbb{E}[|\log{(1+U_{k})}|^{1+\delta}\mid\mathcal{A}^{c}]]$
$\leq 2^{\delta}\mathbb{E}[(|U_{k}|+U_{k}^{2}))^{1+\delta}\mid\mathcal{A}]+2^{\delta}|\log{np_{k}}|^{1+\delta}\cdot\mathbb{P}(\mathcal{A}^{c})$
$\leq C\sum_{k}p_{k}(np_{k})^{-(1+\delta)/2}+2^{1+\delta}|\log{np_{k}}|^{(1+\delta)}e^{-cnp_{k}}$
$\leq Cn^{-(1+\delta)/2}\ell^{(1+\delta)}\Big(\sum_{k}p_{k}\Big)^{1-(1+\delta)/2}+\cdots$
$\leq C(\ell^{2}/n)^{(1+\delta)/2}+2^{1+\delta}|\log{np_{k}}|^{(1+\delta)}e^{-cnp_{k}}$
As $\ell^{2}=o(n)$, the first term goes to 0. Also the second term goes to 0. Thus, $\sup_{n}\mathbb{E}[|\log(n_{k}/np_{k})|^{1+\delta}]<\infty$.
Finiteness of $\mathbb{E}[|\log S_{x}|^{1+\delta}]$ and $\mathbb{E}[|\log S_{y}|^{1+\delta}]$.
We prove this for the $x$ dimension; the proof for the $y$ dimension is identical. Let $P_{x}^{k}=F_{X|C_{k}}(x_{(i+m_{k})}^{k})-F_{X|C_{k}}(x_{(i-m_{k})}^{k})$. This random variable follows a Beta distribution, $P_{x}^{k}\sim\mathrm{Beta}(2m_{k},n_{k}-2m_{k}+1)$. The conditional mean and variance of its logarithm are given in Johnson:
$\mathbb{E}[\log P_{x}^{k}\mid n_{k}]=\psi(2m_{k})-\psi(n_{k}+1),\quad\mathrm{Var}[\log P_{x}^{k}\mid n_{k}]=\psi_{1}(2m_{k})-\psi_{1}(n_{k}+1)$
where $\psi(\cdot)$ and $\psi_{1}(\cdot)$ are the digamma and trigamma functions.
Using the standard asymptotic expansions from NIST, $\psi(z)=\log{z}-\frac{1}{2z}+O(z^{-2})$ and $\log(1+\frac{1}{z})=\frac{1}{z}-\frac{1}{2z^{2}}+O(z^{-3})$, we obtain
$\mathbb{E}[\log P_{x}^{k}\mid n_{k}]=\log\left(\frac{2m_{k}}{n_{k}}\right)+O\left(\frac{1}{m_{k}}\right),\quad\mathrm{Var}[\log{P_{x}^{k}}\mid n_{k}]=\tfrac{1}{2m_{k}}+O\!\Big(\tfrac{1}{m_{k}^{2}}\Big)$
Now we bound the conditional expectation of $|\log S_{x}|^{1+\delta}$:
$\mathbb{E}[|\log S_{x}|^{1+\delta}\mid n_{k}]$
$\leq 2^{\delta}\left|\log\left(\frac{2m_{k}}{n_{k}}\right)-\mathbb{E}[\log P_{x}^{k}\mid n_{k}]\right|^{1+\delta}+2^{\delta}\mathbb{E}\left[\left|\log P_{x}^{k}-\mathbb{E}[\log P_{x}^{k}\mid n_{k}]\right|^{1+\delta}\mid n_{k}\right]$
$\leq 2^{\delta}\left|O\left(\frac{1}{m_{k}}\right)\right|^{1+\delta}+2^{\delta}\left(\mathrm{Var}\left(\log P_{x}^{k}\mid n_{k}\right)\right)^{\frac{1+\delta}{2}}$
$\leq O\left(\frac{1}{m_{k}^{1+\delta}}\right)+2^{\delta}\left(\tfrac{1}{2m_{k}}+O\!\Big(\tfrac{1}{m_{k}^{2}}\Big)\right)^{\frac{1+\delta}{2}}\leq Cn_{k}^{-(1+\delta)/4},$
using $m_{k}\asymp\sqrt{n_{k}}$ and Lyapunov’s Inequality. Now set an event $A_{k}:=\{n_{k}\ \geq\ \tfrac{1}{2}\,np_{k}\}$. By a Chernoff bound, $\mathbb{P}(A_{k}^{c}\mid C_{k})\leq\exp(-c\,np_{k})$. On $A_{k}$, $n_{k}^{-(1+\delta)/4}\leq(2/(np_{k}))^{r/4}$. Therefore,
$\mathbb{E}\big[|\log S_{x}|^{(1+\delta)}\big]=\sum_{k}p_{k}\,\mathbb{E}\big[|\log S_{x}|^{(1+\delta)}\mid n_{k}\big]\leq C\sum_{k}p_{k}(np_{k})^{-(1+\delta)/4}+C\sum_{k}p_{k}e^{-c\,np_{k}}.$
Using Jensen inequality and $\ell^{2}=o(n)$, we show the first term goes to 0.
$n^{-(1+\delta)/4}\sum_{k}p_{k}^{1-(1+\delta)/4}\leq n^{-(1+\delta)/4}(\ell^{2})^{(1+\delta)/4}\Big(\sum_{k=1}^{\ell^{2}}p_{k}\Big)^{1-(1+\delta)/4}\leq(\ell^{2}/n)^{(1+\delta)/4}$
Also $\sum_{k}p_{k}e^{-c\,np_{k}}\leq 1$ for all $n$ (and $\to 0$ if $n\min_{k}p_{k}\to\infty$). Hence, $\sup_{n}\mathbb{E}|\log S_{x}|^{1+\delta}<\infty$. The same argument applies to $\log S_{y}$.
Finiteness of $\mathbb{E}[|\log f(X^{\prime},y_{k}^{*})|^{1+\delta}]$, $\mathbb{E}[|\log f(x_{k}^{*},Y^{\prime})|^{1+\delta}]$, and $\mathbb{E}\big[\big|\log(|C_{k}|/p_{k})\big|^{1+\delta}\big]$
Let $g(u):=|\log f(u)|^{1+\delta}$ with $\int gf<\infty$. Define the upper step function $s_{n}(u):=\sum_{k}\big(\sup_{v\in C_{k}}g(v)\big)\mathbf{1}_{C_{k}}(u)$. Since the mesh $\to 0$ and $f$ is continuous with $f>0$ a.e., $s_{n}\downarrow g$ pointwise and $\int s_{n}f\downarrow\int gf<\infty$. Because $(X^{\prime},y_{k}^{*})\in C_{k}$ and
$\bar{f}_{k}:=|C_{k}|^{-1}\!\int_{C_{k}}f\in[\inf_{C_{k}}f,\sup_{C_{k}}f]$ we have,
$|\log f(X^{\prime},y_{k}^{*})|^{r}\leq\sup_{u\in C_{k}}g(u),\qquad\Big|\log\frac{|C_{k}|}{p_{k}}\Big|^{r}=\big|\log\bar{f}_{k}\big|^{r}\leq\sup_{u\in C_{k}}g(u).$
Hence, $\sup_{n}$ of each is finite; the same for $\mathbb{E}|\log f(x_{k}^{*},Y^{\prime})|^{1+\delta}$.
Since each term in the inequality is uniformly bounded in $L^{1+\delta}$, their sum is also uniformly bounded. Therefore, we have the desired uniform moment control:
$\sup_{n}\mathbb{E}\left[\,|\log\hat{f}_{n,\ell}(X,Y)|^{1+\delta}\,\right]<\infty$. Since $1+\delta>1$, de la Vallée-Poussin’s criterion implies that the sequence $\{\log\hat{f}_{n,\ell}\}$ is uniformly integrable. This, combined with the almost sure convergence $\log\hat{f}_{n,\ell}\xrightarrow{a.s.}\log f$, allows us to apply Vitali’s Convergence Theorem to conclude that $\lim_{n\to\infty}\mathbb{E}[\log\hat{f}_{n,\ell}]=\mathbb{E}[\log f]$.
This proves that the bias term $B_{n}:=\mathbb{E}[\log\hat{f}_{n,\ell}]-\mathbb{E}[\log f]$ converges to 0. The variance term $A_{n}:=-\frac{1}{n}\sum(\log\hat{f}_{n,\ell}(X_{i},Y_{i})-\mathbb{E}[\log\hat{f}_{n,\ell}])$ converges to 0 almost surely by the Strong Law of Large Numbers. Thus, $\hat{H}_{n,\ell}\xrightarrow{a.s.}H(f)$.
To prove $L^{1}$ convergence, we must show the sequence $\{\hat{H}_{n,\ell}\}$ is uniformly integrable. By Jensen’s inequality for the convex function $\phi(z)=|z|^{1+\delta}$:
$\mathbb{E}\bigl[|\hat{H}_{n,\ell}|^{1+\delta}\bigr]\leq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\bigl[|\log\hat{f}_{n,\ell}(X_{i})|^{1+\delta}\bigr]=\mathbb{E}\bigl[|\log\hat{f}_{n,\ell}(X_{1},Y_{1})|^{1+\delta}\bigr]$
Since we showed the right-hand side is uniformly bounded in $n$, we have
$\sup_{n}\mathbb{E}\bigl[|\hat{H}_{n,\ell}|^{1+\delta}\bigr]<\infty$. This implies $\{\hat{H}_{n,\ell}\}$ is uniformly integrable. As $\hat{H}_{n,\ell}\xrightarrow{a.s.}H(f)$, Vitali’s Convergence Theorem again applies, yielding the final result: $\hat{H}_{n,\ell}\xrightarrow{L^{1}}H(f)$.
8.8 Proof Sketch of Theorem 4
The proof follows the same three-step logic as the bivariate proof detailed in Appendix E. First, by the Strong Law of Large Numbers, the sample proportion $\frac{n_{k}}{n}$ converges almost surely to the true partition probability. Second, by Lemma 2, each of the $d$ univariate conditional density estimators, $\hat{f}_{X_{j};k}(x_{j})$, converges almost surely to the true conditional density, $f_{X_{j}|C_{k}}(x_{j})$.
The final step is to show that the limit of the full estimator, $\frac{n_{k}}{n}\prod_{j=1}^{d}\hat{f}_{X_{j};k}(x_{j})$, converges to $f(\mathbf{x})$. By extending the Mean Value Theorem argument from the bivariate case, the product of the limits can be shown to converge to the true joint density.
This establishes the almost sure convergence, $\hat{f}_{n,l}(\mathbf{x})\xrightarrow{a.s.}f(\mathbf{x})$. The $L^{1}$ convergence then follows directly from this result and Proposition 3 by an application of Scheffé’s Theorem.
8.9 Proof Sketch of Theorem 5
The proof establishes the uniform integrability of $\{\log\hat{f}_{n,\ell}(\mathbf{X})\}$ by proving it is bounded in $L^{1+\delta}$. The log-density estimator decomposes as $\log\hat{f}_{n,\ell}(\mathbf{x})=\log(n_{k}/n)+\sum_{j=1}^{d}(\log S_{j}+\log T_{j})$, where
$S_{j}$ $=\frac{2m_{k}}{n_{k}\{F_{X_{j}|C_{k}}(x_{j,(a_{j}+m_{k})}^{k})-F_{X_{j}|C_{k}}(x_{j,(a_{j}-m_{k})}^{k})\}},$
$T_{j}$ $=\frac{F_{X_{j}|C_{k}}(x_{j,(a_{j}+m_{k})}^{k})-F_{X_{j}|C_{k}}(x_{j,(a_{j}-m_{k})}^{k})}{x_{j,(a_{j}+m_{k})}^{k}-x_{j,(a_{j}-m_{k})}^{k}}.$
Applying the MVT along the other (d1)(d-1) coordinates, $T_{j}=\frac{|I_{-j}^{(k)}|}{p_{k}}f(\tilde{X}_{j})$ for some $\tilde{X}_{j}\in C_{k}$. Taking logs and using $\sum_{j=1}^{d}\log|I_{-j}^{(k)}|=(d-1)\log|C_{k}|$ yields the final decomposition:
$\boxed{\ \log\hat{f}_{n,\ell}(X)=\underbrace{\log\tfrac{n_{k}}{np_{k}}}_{\text{(A)}}+\underbrace{\sum_{j=1}^{d}\log S_{j}}_{\text{(B)}}+\underbrace{\sum_{j=1}^{d}\log f(\tilde{X}_{j})}_{\text{(C)}}+\underbrace{(d-1)\log\tfrac{|C_{k}|}{p_{k}}}_{\text{(D)}}\ }$
Let $r=1+\delta$ then, the $L^{r}$ norms of these terms are bounded as follows, where $s_{n}$ is the upper step function of $|\log f|^{r}$:
$\text{(A)}\quad\mathbb{E}|\log\tfrac{n_{k}}{np_{k}}|^{r}$ $\leq C(\tfrac{\ell^{d}}{n})^{r/2}+C,$ $\text{(C)}\quad\mathbb{E}[\sum_{j=1}^{d}|\log f(\tilde{X}_{j})|^{r}]$ $\leq d\int s_{n}f,$
$\text{(B)}\sum_{j=1}^{d}\mathbb{E}|\log S_{j}|^{r}$ $\leq C(\tfrac{\ell^{d}}{n})^{r/4}+C,$ $\text{(D)}\quad\mathbb{E}|(d-1)\log\tfrac{|C_{k}|}{p_{k}}|^{r}$ $\leq(d-1)^{r}\!\int s_{n}f.$
where $\int s_{n}f\downarrow d\int|\log f|^{r}f<\infty$. Combining the bounds gives
$\sup_{n}\mathbb{E}|\log\hat{f}_{n,\ell}(X)|^{r}<\infty$, hence uniform integrability. Almost sure convergence and UI imply $\mathbb{E}[\log\hat{f}_{n,\ell}(X)]\to\mathbb{E}[\log f(X)]$ by Vitali’s theorem. For $\hat{H}_{n,\ell}$, a bounded-difference argument controls the empirical fluctuation, and Jensen’s inequality implies $\sup_{n}\mathbb{E}|\hat{H}_{n,\ell}|^{r}\leq\sup_{n}\mathbb{E}|\log\hat{f}_{n,\ell}(X)|^{r}<\infty$, so $\{\hat{H}_{n,\ell}\}$ is uniformly integrable. Therefore $\hat{H}_{n,\ell}\to H(f)$ a.s. and in $L^{1}$.